OpenStudy (hoybaby):

how do i find the inverse of h=radical 3s over 2????

OpenStudy (hoybaby):

\[h=\sqrt{3s}\over 2\]

OpenStudy (anonymous):

\[ 2/\sqrt{3s}\]

OpenStudy (hoybaby):

what? how did you get that?

OpenStudy (anonymous):

that's if your question is about the multiplication inverse of h

OpenStudy (hoybaby):

the two is not diving the h just the \[\sqrt{3}\]

OpenStudy (hoybaby):

i have to write the formla so that the height is the input and the legth of the sides is a output which means the inverse

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[=\sqrt{3s}/2 ====> 2h=\sqrt{3s} ===> 4h^{2} =3s ====> s =4h^{2}/3\]

OpenStudy (hoybaby):

where do you get the 4 from?

OpenStudy (anonymous):

\[s=4h^{2} /3\]

OpenStudy (anonymous):

squaring both sides will give \[3s =4h^{2}\]

OpenStudy (anonymous):

then divide both sides on 3 \[s=4h^{2}/3\]

OpenStudy (anonymous):

the idea is to take the s out of the radical by squraring both side

OpenStudy (hoybaby):

ok i think i see it much clearer now. care to help me with other problem i have?

OpenStudy (anonymous):

ok

OpenStudy (hoybaby):

\[h=0.8(200-A)\] uf biwlers average is over 200 the handicap is 0. rewrite it so average is output and handicap in input. if your handicap is 32, whats your average?

OpenStudy (anonymous):

\[h=0.8(200-A)\] \[h =160 -0.8 A\] \[0.8 A=160 - h\] \[A=(160-h)/0.8\] when h=32 ===> A =(160-32)/0.8 = 160

OpenStudy (hoybaby):

if my handicap is 32, will that make my average be 134.4?

OpenStudy (anonymous):

let me check again

OpenStudy (anonymous):

I got 160